40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 5 05:16:12 2009, Yoichi, Update, Locking, DC Readout and DARM response DARM-TF.png
    Reply  Tue May 5 08:26:56 2009, rob, Update, Locking, DC Readout and DARM response 
       Reply  Tue May 5 11:44:33 2009, rob, Update, Locking, DARM response DARMtf_nospring.pngDARMtf_demodphases.png
Message ID: 1545     Entry time: Tue May 5 08:26:56 2009     In reply to: 1544     Reply to this: 1548
Author: rob 
Type: Update 
Category: Locking 
Subject: DC Readout and DARM response 

Quote:
Tonight, I was able to switch the DARM to DC readout a couple of times.
But the lock was not as stable as the RF DARM. It lost lock when I tried to measure the DARM loop gain.

I also measured DARM response when DARM is on RF.
The attached plot shows the DARM optical gain (from the mirror displacement to the PD output).
The magnitude is in an arbitrary unit.

I measured a transfer function from DARM excitation to the DARM error signal. Then I corrected it for the DARM open loop gain and the pendulum response to get the plot below.

There is an RSE peak at 4kHz as expected. The origin of the small bump and dip around 2.5kHz and 1.5kHz are unknown.
I will consult with the Optickle model.
I don't know why the optical gain decreases below 50Hz (I don't think it actually decreases).
Seems like the DARM loop gain measured at those frequencies are too low.
I will retry the measurement.


The optical gain does decrease below ~50Hz--that's the optical spring in action. The squiggles are funny. Last time we did this we measured the single arm TFs to compensate for any tough-to-model squiggles in the transfer functions which might arise from electronics or the suspensions.
ELOG V3.1.3-